# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v2

@@ 7.2 Application of GGDs in Graph Classification Tasks
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell4-122 = c, cell4-123 = c, cell4-124 = c, cell4-125 = c, hline1,3,13 = -, hline2 = 2-5, Dataset & Accuracy in percentage
MUTAG PC-3H SW-620H BZR
GGD 86$\pm$7.5 78.34 77.6$\pm$3.5 83.23
TMD, L = 2 76$\pm$5.3
TMD, L = 3 77$\pm$5.2 71.24 70.2$\pm$2.3 73.43
TMD, L = 4 78.2$\pm$6 71.37 70.8$\pm$2.3 73.96
TMD, L = 5 71.89 71.2$\pm$1.88 75.13
GCN(Kipf & Welling, 2017) 77$\pm$3.8 70.56 69.4 72.56
GIN(Xu et al., 2019) 82.6$\pm$4.6 75.34 73.4 77.09
WWL(Togninalli et al., 2019) 72.4$\pm$2.6 65.46 68.34 73.59
WL Subtree(Shervashidze et al., 2011) 76$\pm$6.3 68.43 70.56 N/A
FGH(Titouan et al., 2019) 88.33$\pm$5.6 61.77 59.3 53.66
We evaluate whether the GGD metric aligns with graph labels in graph classification tasks using datasets from TUDatasets (Morris et al., 2020). We employ a Support Vector Classifier (SVC) $(C=1)$ with an indefinite kernel $e^{-\gamma*GGD(G_{1},G_{2})}$, which can be viewed as a noisy observation of the true positive semidefinite kernel (Luss & d’Aspremont, 2007). The parameter $\gamma$ is selected through cross-validation from the set {0.01, 0.05, 0.1}. For comparative analysis with existing methods, we include graph kernels based on graph subtrees: the WL subtree kernel (Shervashidze et al., 2011); and two widely adopted GNNs: graph isomorphism network (GIN) (Xu et al., 2019) and graph convolutional networks (GCN) (Kipf & Welling, 2017).
@@ 7.3 Runtime Complexity Analysis and Comparison
Table 3: Runtime comparison for different distance metrics on graph datasets.
@@
TMD, L = 3 5.30 s 88.6 s 98.7 s 7.22 s
TMD, L = 4 7.89 s 112 s 134.38 s 10.34 s
TMD, L = 6 11.27 s 273 s 288 s 14.98 s
When comparing various graph distance metrics, a primary consideration is their computational complexity. Conventional approaches usually require intricate computations that frequently have cubic time or higher complexities. For our problem, the spectral graph matching step requires the eigenvalue decomposition of adjacency matrices and solving the linear assignment problem (LAP). Eigenvalue decomposition of an $n\times n$ matrix has a complexity of $O(n^{3})$ (Borodin & Munro, 1975; Flamary et al., 2021), while solving the LAP using the Hungarian algorithm also has a runtime complexity of $O(n^{3})$. Similarly, calculating the generalized eigenvalue of two SPD matrices entails a cubic complexity. Consequently, the overall complexity of GGD calculation is $O(n^{3})$. On the other hand, TMD is an OT-based distance metric with a complexity of $O(n^{3}\log(n))$ (Chuang & Jegelka, 2022; Flamary et al., 2021). Therefore, GGD exhibits slightly better (lower) runtime complexity than the TMD metric.
@@
To evaluate runtime performance, we conduct extensive experiments to compare the runtime of computing TMD at various levels with GGD on both small graphs (MUTAG, BZR) and large graphs (PC-3H, SW-620H) collected from the TUDataset (Morris et al., 2020). Table 3 presents the average runtime (in seconds) for computing $100$ distances between different graphs obtained by repeating the experiment five times. The results demonstrate that GGD consistently outperforms TMD in terms of runtime across all datasets and scenarios, particularly when dealing with larger graphs that contain more nodes. The TMD metric computation usually requires more levels to effectively capture the entire graph structure. In such cases, GGD exhibits runtime performance approximately $6-9$ times faster than of TMD. Hence, we conclude that GGD is significantly more computationally efficient than TMD, especially when working with large graphs. More details about our experimental setup can be found in Appendix A.10.
7.4 Partial Node Features
Cutting-edge graph distance metrics like TMD, rely on node attributes to compute the dissimilarity between graphs, resulting in more accurate outcomes when all attributes are available. However, acquiring datasets with complete node attributes is often unattainable in real-world scenarios, leading to situations where certain features are partially missing. In such scenarios when only partial node features are available, we compare TMD with GGD to better understand their differences. Table 4 shows that the TMD metric outperforms GGD at various levels when node features are fully accessible. However, when node features are randomly removed from the MUTAG dataset, the accuracy of TMD degrades substantially.
{tblr}
cell11 = r=2, cell12 = c=5, hline1 = 1-60.08em, hline2 = 2-6, hline3,6-7 = -, Distance
metric & Node features missing in percentage
0% 20% 50% 80% 100%
TMD, L = 3 0.84 0.78 0.72 0.63 0.61
TMD, L = 4 0.81 0.77 0.62 0.58 0.57
TMD, L = 5 0.80 0.75 0.65 0.58 0.53
GGD 0.78 0.78 0.77 0.77 0.77
7.5 GGD Approximation using Extreme Eigenvalues
Table 5: Performance of GGD using extreme eigenvalues only.
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=3, cell32 = c, cell33 = c, cell34 = c, cell42 = c, cell43 = c, cell44 = c, hline1,3,5 = -, hline2 = 2-4, Task & Number of extreme eigenvalues
2 4 All
Correlation with GNN 0.74 0.76 0.77
Classification accuracy $81.50\pm 6.85$ $83.87\pm 7.56$ $86.00\pm 7.5$
The largest and the smallest eigenvalues correspond to the most dominant mismatches in graph cuts and effective resistance distances, contributing the most to the total GGD value. Similarly, the second largest and smallest eigenvalues correspond to the next significant mismatched cuts. In our experiment, we obtain approximate GGDs using a few extreme eigenvalue pairs and compare them with the ground truth. Figure 4 illustrates the relative accuracy of the approximate GGDs, in which we observe that the top four pairs of extreme eigenvalues contribute $80\%$ of the total GGD values. In addition, we conduct the SVC classification task and GNN correlation study using GGD with only $2$ and $4$ extreme eigenvalue pairs, respectively, and present the associated findings in Table 5.
@@ 8 Conclusion
In this work, we introduce Graph Geodesic Distance (GGD), a novel spectral graph distance metric based on graph matching and the infimum on a Riemannian manifold. GGD captures the essential structural mismatches of graphs vital for graph classification tasks. Additionally, we show that GGD can serve as an effective metric for analyzing the stability of GNN models and graph classification tasks, achieving superior performance when only partial node features are available.
@@ Appendix A Appendix
A.1 Graph Adjacency and Laplacian Matrices
@@
Let $D$ denote the diagonal matrix where $D(i,i)$ is equal to the (weighted) degree of node $i$. The graph Laplacian matrix is then given by $L=D-A$. The rank of the Laplacian matrix of a graph $G$ is $n-c(G)$, where $n$ is the number of nodes and $c(G)$ is the number of connected components in the graph. For a connected graph, this implies that the rank of the Laplacian matrix is $n-1$, meaning Laplacian matrices are not full-rank (Bondy et al., 1976).
A.2 Riemannain Manifold
@@
A manifold is a type of topological space that resembles Euclidean space in small, local regions around each point. In other words, for every point on a manifold, there is a neighborhood that is similar to a flat multidimensional space. A Riemannian manifold is a type of manifold equipped with a smoothly varying inner product on the tangent spaces at each point. This means that for every point on the manifold, the tangent space has a way of measuring distances and angles, and these measurements change smoothly from point to point (Lee, 2018). In simpler terms, a Riemannian manifold is a smooth, curved space that locally behaves like Euclidean space but has its own geometric properties, such as how distances, angles, and volumes are defined. These properties are determined by a Riemannian metric, which generalizes the concept of measuring lengths and angles in flat space to curved spaces (Lee, 2018).
@@
A Riemannian manifold can have curvature, unlike a flat space. This curvature allows the study of geometric shapes ranging from spheres and cylinders to more abstract surfaces. The Riemannian structure enables us to compute geodesics, volumes, and various types of curvature. This makes Riemannian manifolds fundamental in fields like differential geometry and physics, and increasingly important in data science, where curved spaces are used to model complex datasets (You & Park, 2021).
A.3 Effective Resistance in Graph Theory
@@
Effective resistance, also known as resistance distance, is a concept in graph theory that measures the electrical resistance between two nodes in a weighted or unweighted graph represented as a network of resistors (Ghosh et al., 2008). It draws an analogy between electrical networks and graphs, helping to quantify how easily current can flow between two nodes, where the edges are treated as resistors. The effective resistance between nodes provides insight into the connectivity between the network. This means two nodes with lower effective resistance values have higher connectivity (Ellens et al., 2011).
A.4 Detailed Proofs Showing GGD is a Metric
A.4.1 Identity Property
@@ A.4.3 Symmetry Property
Given that the inverse of a symmetric matrix is also symmetric, and the product of two symmetric matrices is symmetric, it follows that ($\mathcal{L}_{1}^{-1}\mathcal{L}_{2}$) is a symmetric tensor. Furthermore, the eigenvalues of a symmetric matrix and the eigenvalues of its inverse matrix are inversely related.
So, the eigenvalues of ($\mathcal{L}_{2}^{-1}\mathcal{L}_{1}$) will be $\frac{1}{\lambda_{1}},\frac{1}{\lambda_{2}},\frac{1}{\lambda_{3}},\ldots,\frac{1}{\lambda_{n}}$.
@@ A.4.4 Triangle Inequality
Now, The Frobenius norm $\|X\|_{F}$ is the geodesic length at $d(\exp X,I)=\|X\|_{F}$ (Bonnabel & Sepulchre, 2010). Hence at identity, $d(\mathcal{L},I)=\|\log\mathcal{L}\|_{F}$.
From (Bonnabel & Sepulchre, 2010; You & Park, 2021) we get,
@@
A.5 Graph Matching Recovery
@@
For some constants $C,c>0$, if $1/n^{0.1}<\eta<c/\log n$, then with probability at least $1-5n^{-5}$ for large $n$, it can be proved that the diagonal components of $\widehat{X}^{*}$ are dominant by showing (Fan et al., 2019):
@@
The difference between the similarity matrix $X$ in the presence of noise and the noiseless situation is bounded. If $\eta>1/n^{0.1}$, then for a constant $C>0$, with probability at least $1-2n^{-5}$ for large $n$, it can be shown (Fan et al., 2019):
@@
Assuming lemma A.1 and A.2, for some $c,c^{\prime}>0$ sufficiently small, and by setting $\eta<c/\log n$ and $\sigma<c^{\prime}\eta$, the algorithm ensures that the right sides of both equations 21 and 22 are at most $1/(12\eta^{2})$. Then when $\pi^{*}=id$ (the identity permutation), these lemmas combine to imply:
@@
with probability at least $1-n^{-4}$. So, all diagonal entries of $\widehat{X}$ are larger than all off-diagonal entries, thereby achieving exact recovery (Fan et al., 2019).
A.6 Choosing $\epsilon$ for converting Laplacians to SPD matrices
Laplacian matrices are symmetric positive semi-definite (SPSD) matrices. To convert these to symmetric positive definite (SPD) matrices, we added a diagonal matrix with very small values ($\epsilon$). We used $0.0001$ as the small value ($\epsilon$) in our experiments. When working with Laplacian matrices of a weighted or unweighted graph, values significantly smaller than the edge weights of that graph have a minimal effect on the transformation. We conducted additional experiments with different small values and included the results in Tables 6 and 7. In Table 6, we observed that in our specific case with the MUTAG graph dataset (Morris et al., 2020), where all graphs are unweighted, any value less than 0.001 has an almost negligible influence on the performance of the graph classification task. Additionally, when using values equal to or less than 0.001, the GGD value remains almost the same, as shown in Table 7.

### v3

@@ 7.2 Application of GGDs in Graph Classification Tasks
Dataset MUTAG PC-3H SW-620H BZR
[…]
We evaluate whether the GGD metric aligns with graph labels in graph classification tasks using datasets from TUDatasets (morris2020tudataset, ). We employ a Support Vector Classifier (SVC) $(C=1)$ with an indefinite kernel $e^{-\gamma*GGD(G_{1},G_{2})}$, which can be viewed as a noisy observation of the true positive semidefinite kernel (luss2007support, ). The parameter $\gamma$ is selected through cross-validation from the set {0.01, 0.05, 0.1}. For comparative analysis with existing methods, we include graph kernels based on graph subtrees: the WL subtree kernel (shervashidze2011weisfeiler, ); and two widely adopted GNNs: graph isomorphism network (GIN) (xu2018how, ) and graph convolutional networks (GCN) (Kipf:2016tc, ).
@@
7.3 Application of GGDs in Dataset Distance
To extend the application of GGD beyond structured graph datasets, we explore its utility in measuring distances between datasets. In transfer learning, finding the distance between datasets helps to quantify how similar or dissimilar the source and target domains are, guiding the adaptation of knowledge from one domain to another (otdd, ). We calculated this distance by first converting standard datasets into graph representations and then computing distances between these graphs using GGD framework.
To construct a graph from a dataset, we treat data points as nodes and establish edges based on a k-nearest neighbor (k-NN) approach. Nodes representing similar data points are connected, forming an initial dense graph. To further refine the structure, we apply a spectral graph sparsification procedure to reduce the number of edges while preserving key connectivity properties. The process of creating graphs from the dataset is further explained in Appendix A.17.
Once the graphs are obtained, we apply GGD framework to compute pairwise distances. These distances are then compared with established dataset distance measures based on Optimal Transport (otdd, ). We also explore how these distances correlate with dataset Transferability, which is the improvement in model performance when pretrained on a source dataset and fine-tuned on a target dataset. Transferability $\mathcal{T}$ of a source domain $\mathcal{D}_{S}$ to a target domain $\mathcal{D}_{T}$ is defined as the relative decrease in classification error when adapting compared to training only on the target domain (otdd, ).
$\mathcal{T}\left(\mathcal{D}_{S}\rightarrow\mathcal{D}_{T}\right)=100\times\frac{\operatorname{error}\left(\mathcal{D}_{S}\rightarrow\mathcal{D}_{T}\right)-\operatorname{error}\left(\mathcal{D}_{T}\right)}{\operatorname{error}\left(\mathcal{D}_{T}\right)}.$ (14)
Our results indicate that GGD-based dataset distances provide meaningful insights into dataset relationships and transfer learning ability, matching the performance of conventional distance metrics, demonstrated in Figure 5.
Figure 5: Dataset distance vs. Adaptation for *NIST datasets (M: MNIST, K: KMNIST, F: Fashion-MNIST, U: USPS).
@@ 7.4 Runtime Complexity Analysis and Comparison
When comparing various graph distance metrics, a primary consideration is their computational complexity. Conventional approaches usually require intricate computations that frequently have cubic or higher complexities. For our problem, the spectral graph matching step requires the eigenvalue decomposition of adjacency matrices and solving the linear assignment problem (LAP). Eigenvalue decomposition of an $n\times n$ matrix has a complexity of $O(n^{3})$ (borodin1975computational, ; flamary2021pot, ), while solving the LAP using the Hungarian algorithm also has a runtime complexity of $O(n^{3})$. Similarly, calculating the generalized eigenvalue of two SPD matrices entails a cubic complexity. Consequently, the overall complexity of GGD calculation is $O(n^{3})$. On the other hand, TMD is an OT-based distance metric with a complexity of $O(n^{3}\log(n))$ (chuang2022tree, ; flamary2021pot, ). Therefore, GGD exhibits slightly better (lower) runtime complexity than TMD.
Table 3: Runtime comparison for different distance metrics on various datasets.
@@
TMD, L = 3 5.29s 88.60s 98.69s 7.22s
TMD, L = 4 7.90s 112.12s 134.38s 10.34s
TMD, L = 6 11.27s 273.31s 287.92s 14.98s
@@
To evaluate runtime performance, we compare GGD and TMD on both small graphs (MUTAG, BZR) and large graphs (PC-3H, SW-620H) from the TUDataset (morris2020tudataset, ). Table 3 reports the average time (in seconds) to compute $100$ pairwise distances, averaged over five runs. GGD consistently outperforms TMD across all datasets, especially on larger graphs with more nodes. Since TMD requires deeper levels to capture structural information, its runtime increases rapidly. In contrast, GGD is approximately 6-9 times faster, making it significantly more efficient, particularly on large graphs. More details are provided in Appendix A.18 and Appendix A.12.
8 Limitations
Our current pipeline has three primary limitations. First, computationally, the end-to-end complexity is dominated by eigenvalue decomposition and the linear assignment step (both $O(n^{3})$). Although this is comparable to or slightly better than some OT-based alternatives, scaling to very large graphs remains nontrivial. Appendix A.12 shows that using a small fraction of extreme eigenvalues preserves a high correlation with exact GGD and greatly reduces computation time, but this remains an approximation. Second, GGD is most reliable when comparing graphs of comparable size. When sizes differ dramatically, the size matching via graph coarsening introduces additional approximation error, though we expect such extreme mismatches to be rare in practical applications. Finally, like many graph distance metrics, our theory and experiments assume simple, undirected graphs. Extending GGD to directed, attributed, or higher-order graphs is a valuable direction for future work. We include these limitations to clarify the intended scope and reliability of GGD in practice and to guide when approximations or extensions may be required.
@@ 9 Conclusion
In this work, we introduce Graph Geodesic Distance (GGD), a novel spectral graph distance metric based on graph matching and the infimum on a Riemannian manifold. GGD captures key structural mismatches crucial for graph classification tasks. We also demonstrate that GGD is an effective metric for analyzing GNN model stability and graph classification, delivering superior performance even with partial node features. Additionally, GGD can evaluate transferability by calculating distances between datasets.
@@ Appendix A Appendix
A.1 Algorithm Flow
Algorithm 1 GGD: Geodesic Graph Distance
1:Input: Graphs $G_{1}=(V_{1},E_{1},w_{1})$, $G_{2}=(V_{2},E_{2},w_{2})$, tuning parameter $\eta>0$, small diagonal value $0<\epsilon\ll 1$, node feature weight $\alpha$
2:Output: GGD Value
3:Compute the adjacency matrices $A_{1}$, $A_{2}$
4:if $\text{shape}(A_{1})\neq\text{shape}(A_{2})$ then
5:  Assign the larger graph to $G_{1}$, and the smaller graph to $G_{2}$
6:  while $\text{shape}(A_{1})\geq\text{shape}(A_{2})$ do
7:    Compute the effective resistance $R_{\text{eff}}(p,q)$ of each edge $(p,q)\in E_{1}$
8:    Compute the modified effective resistance $R^{*}_{\text{eff}}(p,q)=R_{\text{eff}}(p,q)+\alpha\lVert NF_{p}-NF_{q}\rVert$
9:    Coarsen the edge with the lowest $R^{*}_{\text{eff}}(p,q)$
10:    Update $A_{1}$
11:  end while
12:end if
13:Compute eigenvectors $u_{i}$, $v_{i}$ and eigenvalues $\zeta_{i}$, $\mu_{i}$ of $A_{1}$ and $A_{2}$, respectively
14:Compute the similarity matrix $\hat{X}\in\mathbb{R}^{n\times n}$
15:Solve Linear Assignment Problem to compute the permutation estimate matrix $\hat{\pi}$
16:Update $A_{2}$ by multiplying with $\hat{\pi}$ to get best match with $A_{1}$
17:Derive $L_{1}$ and $L_{2}$ from $A_{1}$ and $A_{2}$
18:Add $\epsilon$ to diagonal values of $L_{1}$ and $L_{2}$
19:Compute GGD value using the generalized eigenvalues
20:return GGD
A.2 Graph Adjacency and Laplacian Matrices
@@
Let $D$ denote the diagonal matrix where $D(i,i)$ is equal to the (weighted) degree of node $i$. The graph Laplacian matrix is then given by L=DAL=D-A. The rank of the Laplacian matrix of a graph $G$ is nc(G)n-c(G), where $n$ is the number of nodes and $c(G)$ is the number of connected components in the graph. For a connected graph, this implies that the rank of the Laplacian matrix is n1n-1, meaning Laplacian matrices are not full-rank [61].
A.3 Riemannian Manifold
@@
A manifold is a type of topological space that resembles Euclidean space in small, local regions around each point. In other words, for every point on a manifold, there is a neighborhood that is similar to a flat multidimensional space. A Riemannian manifold is a type of manifold equipped with a smoothly varying inner product on the tangent spaces at each point. This means that for every point on the manifold, the tangent space has a way of measuring distances and angles, and these measurements change smoothly from point to point [62]. In simpler terms, a Riemannian manifold is a smooth, curved space that locally behaves like Euclidean space but has its own geometric properties, such as how distances, angles, and volumes are defined. These properties are determined by a Riemannian metric, which generalizes the concept of measuring lengths and angles in flat space to curved spaces [62].
@@
A Riemannian manifold can have curvature, unlike a flat space. This curvature allows the study of geometric shapes ranging from spheres and cylinders to more abstract surfaces. The Riemannian structure enables us to compute geodesics, volumes, and various types of curvature. This makes Riemannian manifolds fundamental in fields like differential geometry and physics, and increasingly important in data science, where curved spaces are used to model complex datasets [63].
A.4 Effective Resistance in Graph Theory
@@
Effective resistance, also known as resistance distance, is a concept in spectral graph theory that draws an analogy between electrical networks and graphs, helping to quantify how easily current can flow between two nodes, where the edges are treated as resistors. The effective resistance between nodes provides insight into the connectivity between the network. This means two nodes with lower effective resistance values have higher connectivity [64].
A.5 Scalable estimation of effective resistances
To address the computational complexity associated with directly computing eigenvalues and eigenvectors required for estimating edge effective resistances, we leverage a scalable framework for approximating the eigenvectors of the graph Laplacian matrix using the Krylov subspace [65]. Let $A$ denote the adjacency matrix of a graph $G$, consider its order-$m$ Krylov subspace $\mathbf{K}_{m}(A,x)$ that is a vector space spanned by the vectors computed through power iterations $x,Ax,A^{2}x,\ldots,A^{m-1}x$ [66]. By enforcing orthogonality among the above vectors in the Krylov subspace, a new set of mutually orthogonal vectors of unit lengths can be constructed for approximating the original Laplacian eigenvectors in 13, which are denoted as $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\ldots$, $\tilde{u}_{m}$. To estimate the effective resistance between two nodes $p$ and $q$, we can exploit the approximated eigenvectors:
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L\tilde{u}_{i}},$ (16)
Graph coarsening with node features
$R^{*}_{eff}(p,q)=R_{eff}(p,q)+\alpha\|f_{p}-f_{q}\|,$ (17)
where $f_{p}$ and $f_{q}$ are node feature vectors of nodes $p$ and $q$, respectively, while $\alpha$ is a weighting factor that determines the effect of node feature information in the graph coarsening process. For instance, if the weight is sufficiently large, the modified effective resistance between nodes with different features will always exceed that of nodes with similar features, effectively preventing their coarsening.
A.6 Detailed Proofs Showing GGD is a Metric
A.6.1 Identity Property
@@ A.6.3 Symmetry Property
Suppose $\lambda_{i}$ is an eigenvalue of $\mathcal{L}_{1}^{-1}\mathcal{L}_{2}$ with corresponding eigenvector $v_{i}$, i.e.,
$\mathcal{L}_{1}^{-1}\mathcal{L}_{2}v_{i}=\lambda_{i}v_{i}.$
Multiplying both sides by $\mathcal{L}_{1}$, we get:
$\mathcal{L}_{2}v_{i}=\lambda_{i}\mathcal{L}_{1}v_{i}.$
Now multiply both sides by $\mathcal{L}_{2}^{-1}$:
$v_{i}=\lambda_{i}\mathcal{L}_{2}^{-1}\mathcal{L}_{1}v_{i}.$
Rearranging, we obtain:
$\mathcal{L}_{2}^{-1}\mathcal{L}_{1}v_{i}=\frac{1}{\lambda_{i}}v_{i}.$
So, the eigenvalues of $\mathcal{L}_{2}^{-1}\mathcal{L}_{1}$ are $\frac{1}{\lambda_{1}},\frac{1}{\lambda_{2}},\ldots,\frac{1}{\lambda_{n}}$, with the same eigenvectors.
@@ A.6.4 Triangle Inequality
Now, The Frobenius norm $\|X\|_{F}$ is the geodesic length at $d(\exp X,I)=\|X\|_{F}$ [36]. Hence at identity, $d(\mathcal{L},I)=\|\log\mathcal{L}\|_{F}$.
From [36, 63] we get,
@@
A.7 Graph Matching Recovery
@@
For some constants $C,c>0$, if $1/n^{0.1}<\eta<c/\log n$, then with probability at least $1-5n^{-5}$ for large $n$, it can be proved that the diagonal components of $\widehat{X}^{*}$ are dominant by showing [67]:
@@
The difference between the similarity matrix $X$ in the presence of noise and the noiseless situation is bounded. If $\eta>1/n^{0.1}$, then for a constant $C>0$, with probability at least $1-2n^{-5}$ for large $n$, it can be shown [67]:
@@
Assuming lemma A.1 and A.2, for some $c,c^{\prime}>0$ sufficiently small, and by setting $\eta<c/\log n$ and $\sigma<c^{\prime}\eta$, the algorithm ensures that the right sides of both equations 22 and 23 are at most $1/(12\eta^{2})$. Then when $\pi^{*}=id$ (the identity permutation), these lemmas combine to imply:
@@
with probability at least $1-n^{-4}$. So, all diagonal entries of $\widehat{X}$ are larger than all off-diagonal entries, thereby achieving exact recovery [67].
A.8 Relation between Generalized Eigenvalues with Cut Mismatch
We selected two graphs from the MUTAG dataset and computed their generalized eigenvalues following the procedure for calculating the Generalized Graph Distance (GGD), which involves determining the node-to-node correspondence. Subsequently, we considerd all possible subsets of nodes and evaluate their corresponding cut mismatches. As shown in Figure 6, each generalized eigenvalue is closely associated with a cut mismatch. This empirical observation supports our hypothesis that the GGD between two input graphs is strongly correlated with structural mismatches in graphs.
Figure 6: Generalized eigenvalues and their corresponding cut mismatches
A.9 Choosing $\epsilon$ for converting Laplacians to SPD matrices
Laplacian matrices are symmetric positive semi-definite (SPSD) matrices. To convert these to symmetric positive definite (SPD) matrices, we added a diagonal matrix with very small values ($\epsilon$). We used $0.0001$ as the small value ($\epsilon$) in our experiments. When working with Laplacian matrices of a weighted or unweighted graph, values significantly smaller than the edge weights of that graph have a minimal effect on the transformation. We conducted additional experiments with different small values and included the results in Tables 4 and 5. In Table 4, we observed that in our specific case with the MUTAG graph dataset [35], where all graphs are unweighted, any value less than 0.001 has an almost negligible influence on the performance of the graph classification task. Additionally, when using values equal to or less than 0.001, the GGD value remains almost the same, as shown in Table 5.